# Fourier Multi-Component and Multi-Layer Neural Networks: Unlocking High-Frequency Potential

### v1

@@ Abstract
The two most critical ingredients of a neural network are its structure and the activation function employed, and more importantly, the proper alignment of these two that is conducive to the effective representation and learning in practice. In this work, we introduce a surprisingly effective synergy, termed the Fourier Multi-Component and Multi-Layer Neural Network (FMMNN), and demonstrate its surprising adaptability and efficiency in capturing high-frequency components. First, we theoretically establish that FMMNNs have exponential expressive power in terms of approximation capacity. Next, we analyze the optimization landscape of FMMNNs and show that it is significantly more favorable compared to fully connected neural networks. Finally, systematic and extensive numerical experiments validate our findings, demonstrating that FMMNNs consistently achieve superior accuracy and efficiency across various tasks, particularly impressive when high-frequency components are present. Our code and implementation details are available here.
@@ 1 Introduction
In this work, we further investigate the behavior and potential of MMNNs and demonstrate a surprising discovery that the $\mathtt{sine}$ function serves as an exceptionally effective activation function for MMNNs. Each component in an MMNN is fundamentally a linear combination of parameterized $\mathtt{sine}$ activation functions and can therefore be viewed as a Fourier series, albeit with relatively small frequency parameters. As a result, each component in an MMNN facilitate the efficient and accurate approximation of a smooth function, while multi-layer compositions can effectively produce high-frequency components, enabling the network to capture more complex function structures with efficient training. Moreover, FMMNN can effectively approximate not only the function but also its derivatives, which can be very important in practice. In the case of approximating a non-smooth function, Fourier approximation can be less effective or result in Gibbs phenomenon. At the same time, as demonstrated in our previous study (Zhang et al., 2023) on shallow networks, activation functions with singularities, such as ReLU, can enhance representational capacity. To address this issue, we introduce a ReLU type of singularity by truncating $\mathtt{sine}$, leading to a novel hybrid activation function called the Sine Truncated Unit ($\mathtt{SinTU}$). Each $\mathtt{SinTU}$ has a form of ${\mathtt{SinTU}}_{s}\coloneqq\sin\circ{\mathcal{T}}_{s}$, where ${\mathcal{T}}_{s}(x)\coloneqq\max\{x,\,s\}.$ The parameter $s$ (typically $\leq 0$) controls the occurrence of singularities and the balance of the hybridization. As $s$ decreases, ${\mathtt{SinTU}}_{s}$ increasingly resembles the $\mathtt{sine}$ function, reducing singularity effects. Moreover, $s$ can be treated as a learnable parameter, either individually for each neuron or shared across all neurons. This variant, denoted as ${\mathtt{PSinTU}}$, is an avenue for future exploration, as its detailed analysis falls beyond the scope of this paper.
We demonstrate that integrating the MMNN structure with $\mathtt{sine}$ (or $\mathtt{SinTU}$s) creates a surprising effective synergy, particularly for efficiently capturing high-frequency components.
@@
  • Finally, extensive numerical experiments demonstrate that Fourier MMNNs (FMMNNs), which use sine or $\mathtt{SinTU}$s as activation functions, consistently outperform other models in both accuracy and efficiency, as shown in Table 1. For $f_{1}$, $\mathtt{sine}$-activated MMNNs achieve the best result, aligning with expectations since $f_{1}$ is $C^{\infty}(\mathbb{R})$ (see Figure 2). The accurate approximation of derivatives is particularly noteworthy, given the complexity of $f_{1}$ and the fact that the training process relies solely on function values, without incorporating derivative information. For $f_{2}\in C^{0}(\mathbb{R})\backslash C^{1}(\mathbb{R})$ (see Figure 2), which contains numerous singularities, ${\mathtt{SinTU}}_{-\pi}$ achieves the best accuracy, demonstrating its effectiveness in capturing these singular features. Notably, even in this inherently challenging case, $\mathtt{sine}$-activated MMNNs still achieve results comparable to the best. For $f_{3}\in C^{0}(\mathbb{R})\backslash C^{1}(\mathbb{R})$ (see Figure 2), FCNNs are highly sensitive to training hyperparameters and often fail with small mini-batches. In contrast, FMMNNs remain stable and perform well across different settings. Even with large mini-batches, training FCNNs is time-consuming, yet they still underperform compared to $\mathtt{sine}$-activated MMNNs. For more details on the experiments, including additional tests in two and three dimensions, refer to Section 3.
@@
The paper is structured as follows. Section 2 provides a detailed analysis of the MMNN architecture, demonstrating its strong approximation capabilities when using sine and $\mathtt{SinTU}$s as activation functions. This section also explores key practical aspects, including the cost function landscape and the interaction between sine and MMNNs, highlighting the advantages of keeping weights fixed within activation functions. The section concludes with a discussion of related work. Section 3 presents extensive numerical experiments that support our theoretical findings. Section 4 provides rigorous proofs for the theorems introduced in Section 2, supplemented by several propositions, whose proofs are detailed in Section 5. Finally, Section 6 concludes the paper with a brief discussion.
@@ 2 The Potential of the $\mathtt{Sine}$ Activation Function
In this section, we first provide a detailed analysis of the MMNN architecture in Section 2.1, followed by an exploration of its mathematical approximation potential when using sine and $\mathtt{SinTU}$s as activation functions in Section 2.2. In Section 2.3, we examine key practical considerations, including the cost function landscape and the interaction between sine and MMNNs, highlighting the benefits of keeping weights within activation functions fixed. The section concludes with a discussion of related work in Section 2.4.
@@ 2.4 Related Work
Some previous works have explored the use of $\mathtt{sine}$ as an activation function in multi-layer networks (Cai et al., 2020; Novello et al., 2024; Morsali et al., 2025; Fathony et al., 2021; Jiao et al., 2023; Sitzmann et al., 2020). The authors in (Novello et al., 2024; Morsali et al., 2025) present numerical examples demonstrating the advantages of sine-activated networks, primarily in FCNNs. Additionally, (Fathony et al., 2021) introduces a multiplicative filter network that incorporates sinusoidal filters and illustrates its benefits through various examples. To the best of our knowledge, existing works do not provide strong mathematical or numerical justification for the benefits of sine, which is a key reason why it is rarely used in practice. This paper first establishes rigorous approximation results, followed by an intuitive landscape analysis of the cost function to assess learning potential, and finally validates these findings through extensive numerical experiments. Our experiments show that using the sine activation in FCNNs does not always lead to good performance. While it works well in some cases, it performs quite poorly in others. We believe this inconsistency is the main reason why sine is not widely adopted in practice. Surprisingly, we found that our MMNN network structure and sine (or $\mathtt{SinTU}$s) form a highly effective combination. Compared to FCNNs, MMNNs exhibit a simpler optimization landscape. Our experiments confirm this observation, as MMNNs achieve consistently strong performance across all test cases.
@@ 3 Numerical Experiments
Section 3.1 compares MMNNs and FCNNs across various activation functions, including $\mathtt{ReLU}$, $\mathtt{GELU}$, tanh, sine, $\mathtt{cosine}$, and $\mathtt{SinTU}$s. Section 3.2 focuses on MMNNs, analyzing the performance of sine relative to other activation functions.
@@ 3.1 MMNNs Versus FCNNs
We employ various network architectures to approximate these functions and evaluate their performance. Notably, these tests are conducted using double precision rather than the default single precision to ensure precise comparisons. For the test corresponding to $f_{1}$, we use a total of 3000 uniformly sampled points from $[-1,1]$ for training. The mini-batch size is set to 3000, meaning all samples are trained simultaneously. The learning rate is defined as $10^{-3}\times 0.9^{\lfloor k/3000\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},300000$ denotes the epoch number. For the test error evaluation, we use another set of 3,000 samples drawn from the uniform distribution $\mathcal{U}(-1,1)$. We emphasize that the mini-batch method is not used for $f_{1}$ because our tests indicate that while mini-batching preserves the approximation of the original function, it leads to poor derivative approximation by only including function values in the cost function. As we can see from Table 2, $\mathtt{sine}$-activated MMNNs perform best as the target function $f_{1}$ is smooth. We point out that the errors for derivatives in Table 2 are relative errors, as absolute errors for derivatives can be misleading (the $L^{\infty}$-norm of $f_{1}^{\prime\prime}$ exceeds 70,000). The accurate approximation of derivatives is surprising, given the complexity of the target function and the fact that the cost function is formulated solely based on the function values, meaning that no information about $f_{1}^{\prime}$ or $f_{1}^{\prime\prime}$ is incorporated into the optimization process.
@@ 3.2 MMNNs: Sine Versus Other Activation Functions
As shown in Table 3, sine and $\mathtt{SinTU}$s are the most effective activation functions for MMNNs. Our results further confirm that the combination of sinusoidal activations and MMNN structures is particularly well-suited for function approximation.

### v2

@@ Abstract
The architecture of a neural network and the selection of its activation function are both fundamental to its performance. Equally vital is ensuring these two elements are well-matched, as their alignment is key to achieving effective representation and learning. In this paper, we introduce the Fourier Multi-Component and Multi-Layer Neural Network (FMMNN), a novel model that creates a strong synergy between them. We demonstrate that FMMNNs are highly effective and flexible in modeling high-frequency components. Our theoretical results demonstrate that FMMNNs have exponential expressive power for function approximation. We also analyze the optimization landscape of FMMNNs and find it to be much more favorable than that of standard fully connected neural networks, especially when dealing with high-frequency features. In addition, we propose a scaled random initialization method for the first layer’s weights in FMMNNs, which significantly speeds up training and enhances overall performance. Extensive numerical experiments support our theoretical insights, showing that FMMNNs consistently outperform traditional approaches in accuracy and efficiency across various tasks. Our code and implementation details are available here.
@@ 1 Introduction
In this work, we further investigate the behavior and potential of MMNNs and demonstrate a surprising discovery that the $\mathtt{sine}$ function serves as an exceptionally effective activation function for MMNNs. Each component in an MMNN is fundamentally a linear combination of parameterized $\mathtt{sine}$ activation functions and can therefore be viewed as a Fourier series, albeit with relatively small frequency parameters. As a result, each component in an MMNN facilitate the efficient and accurate approximation of a smooth function, while multi-layer compositions can effectively produce high-frequency components, enabling the network to capture more complex function structures with efficient training. Moreover, FMMNN can effectively approximate not only the function but also its derivatives, which can be very important in practice, e.g., solving differential equations. Inspired by a scaling strategy for two-layer networks discussed in our previous work (Zhang et al., 2023), which shows that one can reduce the ill-conditioning and improve the representation capacity of the network significantly by scaling the initial weights inside the activation function to be proportional to the network width, we also incorporate the scaling strategy in the initialization of FMMNNs but only to the first layer. Hence, the representation capacity of the first layer is significantly enhanced and can capture more fine features in the samples. This will accelerate the learning process and improve overall performance markedly especially when the target function contains significant high frequency components and ample training samples are available. However, scaling later layers may lead to instability due to fast amplification effect. We present a range of experiments in Section 3.3 to demonstrate the effectiveness of this initial scaling strategy.
In the case of approximating a non-smooth function, Fourier approximation can be less effective or result in Gibbs phenomenon. At the same time, as demonstrated in our previous study (Zhang et al., 2023) on shallow networks, activation functions with singularities, such as ReLU, can enhance representational capacity. To address this issue, we introduce a ReLU type of singularity by truncating $\mathtt{sine}$, leading to a novel hybrid activation function called the Sine Truncated Unit ($\mathtt{SinTU}$). Each $\mathtt{SinTU}$ has a form of ${\mathtt{SinTU}}_{s}\coloneqq\sin\circ{\mathcal{T}}_{s}$, where ${\mathcal{T}}_{s}(x)\coloneqq\max\{x,\,s\}.$ The parameter $s$ (typically $\leq 0$) controls the occurrence of singularities and the balance of the hybridization. As $s$ decreases, ${\mathtt{SinTU}}_{s}$ increasingly resembles the $\mathtt{sine}$ function, reducing singularity effects. Moreover, $s$ can be treated as a learnable parameter, either individually for each neuron or shared across all neurons. This variant, denoted as ${\mathtt{PSinTU}}$, is an avenue for future exploration, as its detailed analysis falls beyond the scope of this paper.
We demonstrate that integrating the MMNN structure with $\mathtt{sine}$ (or $\mathtt{SinTU}$s) creates a surprising effective synergy, particularly for efficiently capturing high-frequency components. Our main contributions are summarized below.
@@
  • Then, extensive numerical experiments demonstrate that Fourier MMNNs (FMMNNs), which use sine or $\mathtt{SinTU}$s as activation functions, consistently outperform other models in both accuracy and efficiency, as shown in Table 1. For $f_{1}$, $\mathtt{sine}$-activated MMNNs achieve the best result, aligning with expectations since $f_{1}\in C^{\infty}(\mathbb{R})$ (see Figure 2). The accurate approximation of derivatives is particularly noteworthy, given the complexity of $f_{1}$ and the fact that the training process relies solely on function values, without incorporating derivative information. Additionally, as we can see from Figure 3, the $\mathtt{sine}$ activation function not only accelerates the convergence of MMNNs but also enhances their overall performance. On the other hand, while ${\mathtt{SinTU}}_{-\pi}$ (with singularity) significantly reduces training error, it results in mediocre test performance. This outcome is expected, as the target function $f_{1}$ is in $C^{\infty}({\mathbb{R}})$. For $f_{2}\in C^{0}(\mathbb{R})\backslash C^{1}(\mathbb{R})$ (see Figure 2), which contains numerous singularities, ${\mathtt{SinTU}}_{-\pi}$ achieves the best accuracy, demonstrating its effectiveness in capturing these singular features. Notably, even in this inherently challenging case, $\mathtt{sine}$-activated MMNNs still achieve results comparable to the best. For $f_{3}\in C^{0}(\mathbb{R})\backslash C^{1}(\mathbb{R})$ (see Figure 2), FCNNs are highly sensitive to training hyperparameters and often fail with small mini-batches. In contrast, FMMNNs remain stable and perform well across different settings. Even with large mini-batches, training FCNNs is time-consuming, yet they still underperform compared to $\mathtt{sine}$-activated MMNNs. For more details on the experiments, including additional tests in two and three dimensions, refer to Section 3.
  • Finally, inspired by scaled parameter initialization in our previous work (Zhang et al., 2023), we propose a scaled random initialization to the weights of the first layer in FMMNNs, which can significantly accelerate the learning process and improve final performance, particularly when ample training samples are available.
@@
(a) $\mathtt{ReLU}$.
(b) $\mathtt{Sine}$.
(c) ${\mathtt{SinTU}}_{-\pi}$.
Figure 3: Illustrations of training and test errors (base-10 logarithm) versus epoch for $f_{1}$.
@@
The paper is structured as follows. In Section 2, we begin with a detailed analysis of the MMNN architecture and its approximation capabilities using sine and $\mathtt{SinTU}$s activation functions. We also discuss practical considerations, including the cost function landscape and the benefits of fixing weights within activation functions, and conclude with a review of related work. Section 3 presents extensive numerical experiments that support our theoretical findings. Section 4 provides rigorous proofs for the theorems introduced in Section 2, supplemented by several propositions, whose proofs are detailed in Section 5. Finally, Section 6 concludes the paper with a brief discussion.
@@ 2 The Potential of the $\mathtt{Sine}$ Activation Function
In this section, we begin with a detailed analysis of the MMNN architecture in Section 2.1, followed by an examination of its mathematical approximation capabilities using sine and $\mathtt{SinTU}$s activation functions in Section 2.2. In Section 2.3, we address practical considerations such as the cost function landscape and the interaction between sine activation functions and MMNNs, emphasizing the advantages of fixing weights within activation functions. We conclude with a discussion of related work in Section 2.4.
@@ 2.4 Related Work
Some previous works have explored the use of $\mathtt{sine}$ as activation functions in neural networks (Cai et al., 2020; Novello et al., 2024; Morsali et al., 2025; Fathony et al., 2021; Jiao et al., 2023; Sitzmann et al., 2020; Wang et al., 2021; Tancik et al., 2020). The authors in (Tancik et al., 2020; Wang et al., 2021) demonstrated that incorporating Fourier Features can enhance neural networks’ ability to learn high-frequency components, although their investigations were limited to FCNNs. To the best of our knowledge, existing works do not provide strong mathematical or numerical justification for the benefits of sine, which is a key reason why it is rarely used in practice. This paper first establishes rigorous approximation results, followed by an intuitive landscape analysis of the cost function to assess learning potential, and finally validates these findings through extensive numerical experiments. Our experiments show that using the sine activation in FCNNs does not always lead to good performance. While it works well in some cases, it performs quite poorly in others. We believe this inconsistency is the main reason why sine is not widely adopted in practice. Surprisingly, we found that our MMNN network structure and sine (or $\mathtt{SinTU}$s) form a highly effective combination. Compared to FCNNs, MMNNs exhibit a simpler optimization landscape. Our experiments confirm this observation, as MMNNs achieve consistently strong performance across all test cases.
@@ 3 Numerical Experiments
Section 3.1 presents a comparison between MMNNs and FCNNs across a variety of activation functions, including $\mathtt{ReLU}$, $\mathtt{GELU}$, tanh, sine, $\mathtt{cosine}$, and $\mathtt{SinTU}$s. Section 3.2 then focuses specifically on MMNNs, evaluating the performance of the sine activation relative to other activation functions. Finally, Section 3.3 uses various numerical experiments to demonstrate that an initial scaling strategy can significantly speed up learning and improve performance when sufficient training data are available.
@@ 3.1 MMNNs Versus FCNNs
We employ various network architectures to approximate these functions and evaluate their performance. Notably, these tests are conducted using double precision rather than the default single precision to ensure precise comparisons. For the test corresponding to $f_{1}$, we use a total of 3000 uniformly sampled points from $[-1,1]$ for training. The mini-batch size is set to 3000, meaning all samples are trained simultaneously. The learning rate is defined as $10^{-3}\times 0.9^{\lfloor k/10000\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},1000000$ denotes the epoch number. For the test error evaluation, we use another set of 3,000 samples drawn from the uniform distribution $\mathcal{U}(-1,1)$. We emphasize that the mini-batch method is not used for $f_{1}$ because our tests indicate that while mini-batching preserves the approximation of the original function, it leads to poor derivative approximation by only including function values in the cost function. As we can see from Table 2, $\mathtt{sine}$-activated MMNNs perform best as the target function $f_{1}$ is smooth. We point out that the errors for derivatives in Table 2 are relative errors, as absolute errors for derivatives can be misleading (the $L^{\infty}$-norm of $f_{1}^{\prime\prime}$ exceeds 70,000). The accurate approximation of derivatives is surprising, given the complexity of the target function and the fact that the cost function is formulated solely based on the function values, meaning that no information about $f_{1}^{\prime}$ or $f_{1}^{\prime\prime}$ is incorporated into the optimization process.
@@ 3.2 MMNNs: Sine Versus Other Activation Functions
As shown in Table 3, sine and $\mathtt{SinTU}$s are the most effective activation functions for MMNNs. Our results further confirm that the combination of sinusoidal activation functions and MMNN structures is particularly well-suited for function approximation.
3.3 Training Acceleration via Scaled Weight Initialization
The initialization of neural network parameters plays a crucial role in determining the efficiency and effectiveness of training. Traditional initialization methods focus on maintaining stable gradients but do not explicitly consider the frequency characteristics of the target function. In our previous study on frequency bias of two-layer networks (Zhang et al., 2023), we showed that one can reduce the frequency bias and enhance the representation ability of the network by scaling the initial random slope weights inside the activation function, which is equivalent to the introduction of high frequency components to the network representation initially. We apply this initial scaling strategy to the first layer of FMMNNs so that more fine features in the samples can be captured by the first layer. We keep the standard initialization in all other layers to avoid instability due to fast multilayer amplification in training. Specifically, we first initialize all parameters by Pytorch default, and then scale ${\bm{W}}_{1},\,{\bm{b}}_{1}$ by $\sqrt{d_{0}}(n_{1}/2)^{1/d_{0}}$, where $n_{1}$ is the width of the first layer and $d_{0}$ is the input dimension (all notation here is based on Section 2.1). It was shown (Zhang et al., 2023) that the best scaling strategy is to make the scaling compatible with the network resolution which is proportional to $n_{1}^{1/d_{0}}$.
The motivation for this approach is to address the spectral bias commonly observed in neural networks, where higher frequency components are learned slower. By initializing the network with higher frequency modes, the model is able to capture more complex patterns at an earlier stage, which results in faster convergence and improved overall performance, particularly when there is sufficient training data. However, the introduction of higher frequency modes in the network can also lead to overfitting if the sample size is limited. Therefore, it is crucial to tailor both the initialization strategy and the network size to the amount of available data.
All training hyperparameters for the experiments shown in Figure 10 are consistent with those outlined in Section 3.1, except for minor adjustments to the mini-batch size and learning rate, as detailed below. For each function $f_{i}$ where $i=1,2,3$, the mini-batch size is set to 200. The learning rate is specified as $10^{-3}\times 0.9^{\lfloor k/m\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},100m$ represents the epoch number, and $m$ is set to 600, 30, and 100 for $f_{1}$, $f_{2}$, and $f_{3}$, respectively. as for Figure 11, For Figure 11, all training hyperparameters are identical to those used in Section 3.2. Additionally, all activation functions employed in these experiments are sine.
(a) MMNN of size (434,16,6).
(b) MMNN of size (434,16,6).
(c) MMNN of size (434,16,6).
(d) MMNN of size (900,16,6).
(e) MMNN of size (900,16,6).
(f) MMNN of size (900,16,6).
(g) FCNN of size (120,–,6).
(h) FCNN of size (120,–,6).
(i) FCNN of size (120,–,6).
Figure 10: Visualization of training and test errors (base-10 logarithm) versus epoch for default and scaling initialization. The first, second, and third columns represent $f_{1}$, $f_{2}$, and $f_{3}$ from Section 3.1, respectively.
(a) MMNN of size (1024,16,6).
(b) MMNN of size (1024,36,8).
(c) ResMMNN of size (1024,36,10).
Figure 11: Visualization of training and test errors (base-10 logarithm) versus epoch for default and scaling initialization. (a), (b), and (c) correspond to $f_{1}$, $f_{2}$, and $f_{3}$ from Section 3.2, respectively.
Figures 10, 11, and 12 demonstrate that our scaled weight initialization approach is particularly effective when there are enough training samples, resulting in significantly faster learning and better final performance. When the sample size is small, as shown in the first column of Figure 10 for $f_{1}$ (a highly oscillatory function) with only 3000 samples, the MMNN tends to overfit. This overfitting occurs because the network width is too large, which introduces excessively high frequencies at initialization. In contrast, the FCNN does not exhibit overfitting in this case, since its width is much smaller.
(a) MMNN of size (120,16,6).
(b) MMNN of size (120,24,10).
(c) MMNN of size (120,32,6).
(d) MMNN of size (180,16,6).
(e) MMNN of size (180,24,6).
(f) MMNN of size (180,32,6).
Figure 12: Visualization of training and test errors (base-10 logarithm) versus epoch for $f_{1}$ from Section 3.1 using our scaling initialization.
Therefore, it is important to adjust the initialization method and select an appropriate network size based on the available sample size. In Figure 12, we present additional experiments with the same settings except for a smaller network size. As shown in the figure, overfitting is avoided and the results are quite good. Finally, we note that our scaling initialization is also effective for FCNNs, as demonstrated by the results in the third row of Figure 10. The results clearly show that our scaling initialization brings significant improvements.